OpenStudy (anonymous):

Medal if you can figure it out but you must explain how you got the answer: A train travels 42 kilometers in 15 minutes. How far will the train travel in 3 hours at that speed? A. 168 km B. 336 km C. 504 km D. 672 km

OpenStudy (hoblos):

the train travels 42 km each 15 minutes can you calculate how many km it travels each 1 minute ?

OpenStudy (anonymous):

I got 504 or C. was that correct? can you help me with another question it'll be the last one... Anton's family drove 216 miles to the lake averaging 48 mi/hr. On the return trip home they averaged 54 mi/hr. What was the total time that Anton's family spent driving to and from the lake? A. 8 h B. 8.5 h C. 9 h D. 9.5 h

OpenStudy (hoblos):

for the going: 48 miles takes 1 hr 216 miles takes x hr x = 216*1/48 = 4.5 hr for the return: 54 miles takes 1 hr 216 miles take 216*1/54 = 4 hr so the total time is 9.5 hr

OpenStudy (anonymous):

oh wow thanks you didn't have to solve it for me but okay thank you @hoblos

OpenStudy (hoblos):

any time :)